WHAT IS THE RATIONAL EQUATION IF DAN AND LAURA TAKE 9 HOURS TOGETHER TO MAKE WEDDING INVITATIONS, BUT IT TAKES LAURA 15 HOURS ALONE. HOW LONG WILL IT TAKE DAN BY HIMSELF.
Answer by stanbon(48569)

Let amount of time it takes Dan alone be "x".
His rate is 1/x per hour.
Laura takes 15 hours; her rate is 1/15 per hour.
Together they take 9 hours; their combined rate is 1/9 per hour.
EQUATION:
Dan rate + Laura rate = together rate
1/x + 1/15 = 1/9
Multiply thru by 45x to get:
45+3x = 5x
2x=45
x=22.5 hours.
Dan takes 22.5 hours to do the job alone
Cheers,
Stan H.